# Wasserstein Wormhole: Scalable Optimal Transport Distance with Transformers

### v1

@@ Abstract
Optimal transport (OT) and the related Wasserstein metric ($W$) are powerful and ubiquitous tools for comparing distributions. However, computing pairwise Wasserstein distances rapidly becomes intractable as cohort size grows. An attractive alternative would be to find an embedding space in which pairwise Euclidean distances map to OT distances, akin to standard multidimensional scaling (MDS). We present Wasserstein Wormhole, a transformer-based autoencoder that embeds empirical distributions into a latent space wherein Euclidean distances approximate OT distances. Extending MDS theory, we show that our objective function implies a bound on the error incurred when embedding non-Euclidean distances. Empirically, distances between Wormhole embeddings closely match Wasserstein distances, enabling linear time computation of OT distances. Along with an encoder that maps distributions to embeddings, Wasserstein Wormhole includes a decoder that maps embeddings back to distributions, allowing for operations in the embedding space to generalize to OT spaces, such as Wasserstein barycenter estimation and OT interpolation. By lending scalability and interpretability to OT approaches, Wasserstein Wormhole unlocks new avenues for data analysis in the fields of computational geometry and single-cell biology.
@@ 1 Introduction
Many problems in statistical learning boil down to computing and comparing probability distributions that fit empirical data, and different approaches vary in their attention to distribution shape, spread, and degree of parameterization. Optimal transport methods attempt to find an ideal pairing between individual data points from two distributions [1] by representing distances as the minimum cost of mapping between (weighted) point clouds. While original OT solvers, which rely on linear programming, scale poorly [2], newer approaches can reduce the computational complexity and generalize OT to new domains such as computational biology [3, 4, 5], numerical geometry [6], chemistry [7], and image processing [8].
We are specifically interested in massive cohorts of empirical distributions, each with up to thousands of points. The computational complexity of finding the Wasserstein distance between two point clouds with Sinkhorn’s algorithm [9], the current favored [10] OT solver, is $O(n^{2}\cdot I)$, where $n$ is the size of each point cloud and $I$ is the number of Sinkhorn iterations until convergence. Despite improvements in scalability, current OT algorithms are unable to analyze large cohorts as we must repeat Sinkhorn’s algorithm for all pairs in the cohorts, requiring $O(N^{2}\cdot n^{2}\cdot I)$ time to analyze $N$ distributions.
@@
Point clouds contain a variable number of unordered samples, hence neural architectures that require a fixed-sized or ordered input are not suited to encode them. On the contrary, transformers [11] are uniquely poised to do so, due to their permutation equivariance, indifference to input length, and ability to model relationships between samples in each point cloud via attention.
In addition to an encoder that embeds point clouds, Wormhole trains a decoder that reconstructs point clouds from their embedding. The decoder facilitates interpretability with its ability to explicitly compute Wasserstein barycenters [12] or interpolate between point clouds [13].
@@
The problem solved by Wormhole is analogous to multidimensional scaling [14]. Both methods find an embedding that preserves distances between samples. While several distance metrics are provably capable of being represented in this manner, it has been shown that a host of other ’non-Euclidean’ metrics, including Wasserstein, cannot be exactly equated with Euclidean distance. Counter-intuitively, classical MDS on non-Euclidean distance matrices increases in error as embedding dimensions increase [15].
Inspired by [15], we extend MDS theory to derive upper and lower bounds on the error incurred by our gradient descent-based approach, thereby providing Wormhole with theoretical guarantees on performance for (non-Euclidean) Wasserstein distances. We further describe a projected gradient descent (PGD) [16] algorithm for MDS which is guaranteed to converge to the global optimum for any distance matrix. On small examples where it is feasible to compute bounds and apply PGD, we empirically show that Wormhole approaches the best solution.
@@ 2 Prior Work
Learning from point-cloud data is dynamic area of research within machine learning [17], with many Optimal Transport based application. PointNet [18, 19] introduced the idea of permutation-invariant fully-connected neural networks to classify and segment 3D point-clouds, which was later extended into transformer based networks [20]. In the context of OT, several works have noted the expressiveness of Wasserstein space compared to euclidean space [21, 22, 23], and have developed methods to embed point-clouds into Wasserstein space, from which relevant information can be more reliably extracted. These works are in essence the inverse of Wormhole, as their target space is Wasserstein, where in Wormhole, target space is euclidean.
Wasserstein Wormhole belongs to a class of algorithms that exploit structures within distributions or relationships between cohorts of them to accelerate or approximate OT. Assuming the transport plan between two distribution is structured, Low-Rank (LR) Sinkhorn [24] factorizes the transport matrix, forcing it into a low-rank structure which enables faster computation of Sinkhorn iterations. Similarly, Convolutional and Geodesic Sinkhorn [25, 26] construct solvers which leverage the underlying geometry of the point-cloud space to simplify OT optimization. MetaOT [27] recognizes that convergence of Sinkhorn depends greatly on quality of initialization and learns a neural network which learns an initial Sinkhorn potential for a given pair of input distributions. Due to its fully-connected architecture, MetaOT is constrained to distributions which share supports, and can not be applied to arbitrary point-clouds. Although these methods can speed up the computation of an individual OT problem, none alleviate the need for many expensive pairwise computations, and are far slower than Wormhole on large cohorts, as shown in Figure 2.
Like Wormhole, methods such as Deep Wasserstein Embeddings (DWE) [28] and DiffusionEMD [29] produce embeddings that preserve Wasserstein distances. DWE learns a neural network via a Siamese architecture trained to match distances, and DiffusionEMD extends the wavelet-based approximation of EMD [30] onto graph domains. Unlike our Wormhole approach, which applies to point clouds in any dimension, DWE is constrained to small 2D images or 3D grids, while scalability of DiffusionEMD is limited to distributions that share supports, such as grayscale images or voxelized spaces. We benchmark Wormhole against DiffusionEMD & DWE, and show that our method is superior at matching OT distances and deriving insights from embeddings, and Wormhole is the only embedding method which scales to high-dimensional datasets (Table 1).
@@
Wormhole follows a long line of neural network approaches for Optimal Transport estimation [31]. Sinkhorn iterations, which are quadratic with size of point-clouds, cannot be applied when When point-clouds are massive consisting of thousands of points. Neural OT [32] parameterizes the transport plan between two empirical distributions as a input-convex neural network [33] which allows for fast, mini-batch training. Such models have been incredibly successful at extending OT to scRNA-seq data [34, 5, 35], where samples commonly contain over $10^{5}$ cells. Wormhole works best in a different setting, where individual OT problems can be solved, but their solvers are not efficient enough to scale to the many pairwise comparisons within a large cohort. Future extensions of our work can make use of neural solvers to unlock Wasserstein analysis to large cohorts of massive point-clouds.
@@ 3.1 Optimal Transport
We follow the formulation of optimal transport presented in [2]. Given two discrete distributions $P_{x}=\sum_{i=1}^{n}\mu_{i}\delta_{x_{i}}$ and $P_{y}=\sum_{i=1}^{m}\nu_{i}\delta_{y_{i}}$ ($\delta$ is the Dirac function), their entropically regularized OT distance is:
@@
Where $U(\nu,\mu):=\{P\in R_{+}^{n,m}:\sum_{j}P_{i,j}=\nu_{i},\sum_{i}P_{i,j}=\mu_{j}\}$ is the set of all valid pairing matrices, $c(x_{i},y_{j})$ is the ground cost function between points in $P_{x}$ and $P_{y}$, usually chosen to be the Euclidean distance, and $H(P)=-\sum_{i,j}P_{i,j}(\log P_{i,j}-1)$ is the entropic term. While the unregularized problem incurs a heavy computational burden for distributions with large supports or in high dimensions, we can compute $W_{\varepsilon}$ via a series of GPU-accelerable vector-matrix products known as Sinkhorn iterations [9].
Due to the regularization term, the $W_{\varepsilon}$ distance between a distribution and itself is not 0, which can introduce biases when used to optimize generative models. To allow for effective optimization while retaining the computational advantages of entropic regularization, a common alternative is the Sinkhorn Divergence [36], which subtracts the self-transport costs from $W_{\varepsilon}$:
@@ 3.2 Multi-dimensional Scaling
From [37], a matrix $D$ is Euclidean if and only if all of its values are non-negative, it is hollow (diagonals are 0) and the criterion matrix $F$ is positive-semidefinite (PSD):
@@
Where $J=I_{N}-\frac{\mathbf{1}_{N}\mathbf{1}_{N}^{\top}}{N}$ is the centering matrix and $1_{N}$ is a vector of length $N$ filled with ones. From brevity, we will call matrices which fit these criterion Euclidean.
Wasserstein distance is not Euclidean (see section 8.3 in [2] and [38]), and there are many examples of cohorts of distributions (Figures S1,S2) wherein pairwise OT distance matrices do not fit the criteria in equation 4. Unlike the Euclidean scenario where classical MDS (cMDS) improves as embedding size grows, on non-Euclidean distance matrices the error behaves counter-intuitively and worsens as dimension increases [15]. This motivates us to seek a description of the theory underpinning the embeddings of non-Euclidean distances.
Dataset Parameters OT Correlation Label Accuracy
MNIST 70,000 105 2 0.986 0.845 0.923 0.862 0.98 0.84 0.97 0.49
Fashion-MNIST 70,000 356 2 0.999 0.971 0.985 0.992 0.87 0.77 0.87 0.73
ModelNet40 12,311 2048 3 0.996 0.852 0.979 0.952 0.86 0.69 0.78 0.61
ShapeNet 15,011 2048 3 0.995 0.817 0.977 0.973 0.98 0.94 0.97 0.92
Rotated ShapeNet (GW) 15,011 512 3 0.987 NA NA 0.687 0.82 NA NA 0.16
MERFISH Cell Niches 256,058 11 254 0.977 OOM OOM OOM 0.98 OOM OOM OOM
SeqFISH Cell Niches 57,407 29 351 0.984 OOM OOM OOM 0.97 OOM OOM OOM
scRNA-seq Atlas 2,185 69 2500 0.986 OOM OOM OOM 0.96 OOM OOM OOM
Table 1: Summary of each dataset used in this study and results from every benchmarked algorithm. Cloud size denotes the median number of points in each point cloud for every dataset. Pearson correlations are computed between true and estimated Wasserstein distances according to pairwise matrices of $10$ instances of $128$ randomly chosen point clouds from each dataset (OT Correlation). The mean squared error (MSE) follows the same trend with Wormhole outperforming competing methods (Table 3). Object classification accuracy is based on the embeddings of each algorithm. After computing encodings, a classifier was trained on the training set embeddings and we denote the classifier’s accuracy on the test set (Label Accuracy). Rotated ShapeNet is the ShapeNet dataset where each sample is randomly rotated in 3D, subsampled by a factor of $4$, and embedded via Gromov-Wasserstein (GW). We divide the results into Wasserstein on 2D and 3D point-clouds, Gromov-Wasserstein on rotated data with arbitrary orientation and high-dimensional datasets. Full implementation details are described in the supplement (Section 11). NA, Not Applicable. OOM, Out-Of-Memory.
@@ 4 Wasserstein Wormhole
  Input: point clouds $\Omega=\{X_{i}\}^{N}_{i=1}$, initialized encoder $T$ and decoder $G$, batch size $B$, learning rate $\varepsilon_{t}$
@@
The complexity of modern OT solvers is quadratic with respect to the number of points in a point cloud; thus, it is only straightfoward to compute pairwise Wasserstein distance for small cohorts of empirical distributions. Indeed, naive approaches will not even scale to standard datasets, such as the 12,311-point-cloud ModelNet40 benchmark [39].
  • Point clouds contain varying numbers of samples. $T(X)$ must be able to generalize to arbitrary input length.
By training the encoder network $T(X)$ (Figure 1, Algorithm 1), we can produce encodings which allow for OT-based analysis of point clouds in linear time, enabling methods such as clustering and visualization [25] on the entire cohort. However, the encoder alone does not provide a clear interpretation of these analyses.
@@ 5 Embedding Non-Euclidean Distance Matrices
Table 2: Comparison between embedding algorithms and theoretical bounds on small datasets where full Wasserstein distance matrices are realizable. Values represent the stress between the embedding and true distance matrices and parentheses denote the size of each dataset. Current MDS embedding algorithms like cMDS and non-linear SMACOF produce far from optimal errors, falling outside our bounds. Wormhole training is stochastic, yet it still converges between the error bounds and approaches the error obtained by the theoretically optimal PGD. Both PGD and Wormhole embeddings were well correlated ($\rho=0.999$) with OT distances for all datasets.
Dataset Lower Upper PGD (our) Wormhole (our) cMDS SMACOF
Simplex ($35$) $0.765$ $6.369$ $1.117$ $1.420$ $7.134$ $5.994$
Gaussian ($128$) $0.129$ $2.552$ $0.168$ $0.401$ $2.681$ $11.729$
MNIST ($256$) $0.042$ $0.616$ $0.058$ $0.100$ $0.657$ $3.231$
In the case where $D$ is Euclidean, by definition there exists a set of embeddings in dimension $d=N$ where the distance matrix is perfectly reconstructed and $\mathcal{L}=0$. Wormhole, however, is trained to minimize the stress relative to the Wasserstein distance matrix, which is generally not Euclidean. It is currently unclear to what extent $D$ can be represented with embeddings when it is not Euclidean.
@@
This bound also appears in [15], which analyzed pathologies of cMDS on non-Euclidean distance matrices.
@@
Algorithm 2 Theoretically Optimal Embedding of non-Euclidean Distance Matrices with Projected Gradient Descent
  Input: Distance Matrix $D$, Initial Solution $D^{\prime}_{0}$, Learning Rate $\varepsilon$
@@
     $D^{*}_{t+1}=D^{\prime}_{t}-\varepsilon\nabla(\|D^{\prime}_{t}-D\|_{F}^{2})$GD Step}
@@
This new projection operator, along with the simple projections onto $C_{2}$ (hollow matrices, by subtraction of the diagonal) and $C_{3}$ (non-negative matrices, by thresholding), allows us to formulate a PGD [16, 40] scheme (Algorithm 2) to embed non-Euclidean distance matrices. This algorithm projects onto $C_{1}\cup C_{2}\cup C_{3}$ using Dykstra iterations [40] and is guaranteed to find the optimum for equation 6, unlike other MDS approaches such as cMDS or SMACOF [41]. The PGD algorithm requires knowing the full pairwise distance matrix and is therefore infeasible in our setting of massive cohorts. However, for small cohorts, the theoretical assurances of PGD provide a useful benchmark for Wormhole.
Our analysis indicates that in real-world datasets, the Wormhole approximation is highly accurate and commensurate with theoretical values. Although this assessment is based on a small subsample of the full dataset, we assume that if random subsets of the larger dataset can be well embedded, the embedding of the entire dataset is also reliable (Table 2). Full derivation of the theoretical bounds, and description of the PGD algorithm appear in Supplementary Sections 9 & 10.
Fig. 3: Wormhole on MNIST point clouds. a. Example of a point cloud produced by thresholding MNIST images. b. Retrieval of the Wasserstein distance in Wormhole embedding space. From a random sample of $128$ point clouds from a test set, we denote the correlation between the true pairwise OT distance and the Euclidean distance between their embeddings. The line $y=x$ is drawn in blue for reference. c. 2D UMAP visualization of Wormhole embeddings of the training and test set point clouds, which recapitulates ground-truth image digit without utilizing label information during training.

### v4

@@ Abstract
Optimal transport (OT) and the related Wasserstein metric ($W$) are powerful and ubiquitous tools for comparing distributions. However, computing pairwise Wasserstein distances rapidly becomes intractable as cohort size grows. An attractive alternative would be to find an embedding space in which pairwise Euclidean distances map to OT distances, akin to standard multidimensional scaling (MDS). We present Wasserstein Wormhole, a transformer-based autoencoder that embeds empirical distributions into a latent space wherein Euclidean distances approximate OT distances. Extending MDS theory, we show that our objective function implies a bound on the error incurred when embedding non-Euclidean distances. Empirically, distances between Wormhole embeddings closely match Wasserstein distances, enabling linear time computation of OT distances. Along with an encoder that maps distributions to embeddings, Wasserstein Wormhole includes a decoder that maps embeddings back to distributions, allowing for operations in the embedding space to generalize to OT spaces, such as Wasserstein barycenter estimation and OT interpolation. By lending scalability and interpretability to OT approaches, Wasserstein Wormhole unlocks new avenues for data analysis in the fields of computational geometry and single-cell biology. Software is available at http://wassersteinwormhole.readthedocs.io/en/latest/.
Machine Learning, ICML
@@ 1 Introduction
Many problems in statistical learning boil down to computing and comparing probability distributions that fit empirical data, and different approaches vary in their attention to distribution shape, spread, and degree of parameterization. Optimal transport (OT) methods attempt to find an ideal pairing between individual data points from two distributions (Villani et al., 2009) by representing distances as the minimum cost of mapping between (weighted) point clouds. While original OT solvers, which rely on linear programming, scale poorly (Peyré et al., 2019), newer approaches can reduce the computational complexity and generalize OT to new domains such as computational biology (Nitzan et al., 2019; Schiebinger et al., 2019; Bunne et al., 2023), numerical geometry (Su et al., 2015), chemistry (Wu et al., 2023), and image processing (Feydy et al., 2017).
We are specifically interested in massive cohorts of empirical distributions, each with up to thousands of points. The computational complexity of finding the Wasserstein distance between two point clouds with Sinkhorn’s algorithm (Cuturi, 2013), the current favored (Flamary et al., 2021) OT solver, is $O(n^{2}\cdot I)$, where $n$ is the size of each point cloud and $I$ is the number of Sinkhorn iterations until convergence. As we must repeat Sinkhorn’s algorithm for all pairs in a cohort, requiring $O(N^{2}\cdot n^{2}\cdot I)$ time to analyze $N$ distributions, current OT algorithms are unable to analyze large cohorts despite improvements in scalability.
@@
Point clouds contain a variable number of unordered samples, hence neural architectures that require a fixed-sized or ordered input are not suited to encode them. On the contrary, transformers (Vaswani et al., 2017) are uniquely poised to do so, due to their permutation equivariance, indifference to input length, and ability to model relationships between samples in each point cloud via attention.
@@
In addition to an encoder that embeds point clouds, Wormhole trains a decoder that reconstructs point clouds from their embedding. The decoder facilitates interpretability with its ability to explicitly compute Wasserstein barycenters (Cuturi & Doucet, 2014) or interpolate between point clouds (Chewi et al., 2021).
The problem solved by Wormhole is analogous to multidimensional scaling (Torgerson, 1952). Both methods find an embedding that preserves distances between samples. While several distance metrics are provably capable of being represented in this manner, it has been shown that a host of other ”non-Euclidean” metrics, including Wasserstein, cannot be exactly equated with Euclidean distance. Counter-intuitively, classical MDS on non-Euclidean distance matrices becomes more inaccurate as embedding dimensions increase (Sonthalia et al., 2021).
Inspired by (Sonthalia et al., 2021), we extend MDS theory to derive upper and lower bounds for the error incurred by embedding non-Euclidean distances, such as Wasserstein. We further describe a projected gradient descent (PGD) (Boyd & Vandenberghe, 2004) algorithm for MDS which is guaranteed to converge to the global optimum for any distance matrix. On small examples where it is feasible to compute bounds and apply PGD, we find that Wormhole approaches the optimal solution and falls within our derived bounds.
Dataset Parameters OT Correlation Label Accuracy
MNIST 70,000 105 2 0.98 0.845 0.92 0.86 0.98 0.84 0.97 0.49
Fashion-MNIST 70,000 356 2 0.99 0.97 0.98 0.99 0.87 0.77 0.87 0.73
ModelNet40 12,311 2048 3 0.99 0.85 0.97 0.95 0.86 0.69 0.78 0.61
ShapeNet 15,011 2048 3 0.99 0.81 0.97 0.97 0.98 0.94 0.97 0.92
Rotated ShapeNet (GW) 15,011 512 3 0.98 NA NA 0.68 0.82 NA NA 0.16
MERFISH Cell Niches 256,058 11 254 0.97 OOM OOM OOM 0.98 OOM OOM OOM
SeqFISH Cell Niches 57,407 29 351 0.98 OOM OOM OOM 0.97 OOM OOM OOM
scRNA-seq Atlas 2,185 69 2500 0.98 OOM OOM OOM 0.96 OOM OOM OOM
Table 1: Summary of each dataset used in this study and results from every benchmarked algorithm. Cloud size denotes the median number of points in each point cloud for every dataset. Pearson correlations are computed between true and estimated Wasserstein distances according to pairwise matrices of $10$ instances of $128$ randomly chosen point clouds from each dataset (OT Correlation). The mean squared error (MSE) follows the same trend with Wormhole outperforming competing methods (Table 3). Object classification accuracy is based on the embeddings of each algorithm. After computing encodings, a classifier was trained on the training set embeddings and we denote the classifier’s accuracy on the test set (Label Accuracy). Rotated ShapeNet is a version of the standard ShapeNet dataset where each sample is randomly rotated in 3D, subsampled by a factor of $4$, and embedded via Gromov-Wasserstein (GW). We divide the results into Wasserstein on 2D and 3D point-clouds, Gromov-Wasserstein on rotated data with arbitrary orientation and high-dimensional datasets. Full implementation details are described in the supplement (Section 11). NA, Not Applicable. OOM, Out-Of-Memory.
@@ 2 Prior Work
There have been many prior works on learning from point cloud data (Guo et al., 2020), often involving OT-based solutions. PointNet (Garcia-Garcia et al., 2016; Qi et al., 2017) introduced the idea of permutation-invariant, fully-connected neural networks to classify and segment 3D point-clouds, which was later extended into transformer-based networks (Lee et al., 2019). In the context of OT, several works have noted the expressiveness of Wasserstein space compared to euclidean space (Frogner et al., 2019; Katageri et al., 2024; Liu et al., 2022), and have developed methods to embed point-clouds into Wasserstein space, from which relevant information can be more reliably extracted. These works are in essence the inverse of Wormhole, as their target space is Wasserstein, in contrast to Wormhole’s Euclidean target space.
@@
Wasserstein Wormhole belongs to a class of algorithms that exploit structures within distributions or relationships between cohorts of them to accelerate or approximate OT. Low-Rank (LR) Sinkhorn (Scetbon et al., 2021) factorizes the pairing matrix into a low-rank composition, assuming the transport plan between two distribution is structured, to enable faster computation of Sinkhorn iterations. Convolutional and Geodesic Sinkhorn (Bachmann et al., 2022; Solomon et al., 2015) construct solvers which leverage the underlying geometry of the point-cloud space to simplify OT optimization. MetaOT (Amos et al., 2022) recognizes that convergence of Sinkhorn depends greatly on quality of initialization and fits a neural network which learns an initial Sinkhorn potential for a given pair of input distributions. Due to its fully-connected architecture, MetaOT is constrained to distributions which share supports, and cannot be applied to arbitrary point-clouds. Although these methods can speed up the computation of an individual OT problem, none alleviate the need for many expensive pairwise computations, and are far slower than Wormhole on large cohorts (Figure 2).
Previous works have also employed Wasserstein preserving euclidean embeddings in the pursuit of scaling OT to large cohorts (Tong et al., 2021), some using neural network models (Courty et al., 2017; Sehanobish et al., 2020). DWE (Courty et al., 2017) trains a Siamese neural network to match distances, and DiffusionEMD (Tong et al., 2021) extends the wavelet-based approximation of EMD (Shirdhonkar & Jacobs, 2008) onto graph domains. Unlike our Wormhole approach, which applies to point clouds in any dimension, DWE is constrained to small 2D images or 3D grids, while scalability of DiffusionEMD is limited to distributions that share supports, such as grayscale images or voxelized spaces. We benchmark Wormhole against DiffusionEMD & DWE, and show that our method is superior at matching OT distances and deriving insights from embeddings on low-dimensional point clouds. On general datasets, Wormhole is the only embedding method which can scale to high-dimensions.
@@ 3.1 Optimal Transport
We follow the formulation of optimal transport presented in (Peyré et al., 2019). Given two discrete distributions $P_{x}=\sum_{i=1}^{n}\mu_{i}\delta_{x_{i}}$ and $P_{y}=\sum_{i=1}^{m}\nu_{i}\delta_{y_{i}}$ ($\delta$ is the Dirac function), their entropically regularized OT distance is:
@@
Where $U(\nu,\mu):=\{P\in R_{+}^{n,m}:\sum_{j}P_{i,j}=\nu_{i},\sum_{i}P_{i,j}=\mu_{j}\}$ is the set of all valid pairing matrices, $c(x_{i},y_{j})$ is the ground cost function between points in $P_{x}$ and $P_{y}$, usually chosen to be the Euclidean distance, and $H(P)=-\sum_{i,j}P_{i,j}(\log P_{i,j}-1)$ is the entropic term. While the unregularized problem incurs a heavy computational burden for distributions with large supports or in high dimensions, $W_{\varepsilon}$ can be computed via a series of GPU-accelerable vector-matrix products known as Sinkhorn iterations (Cuturi, 2013).
Due to the regularization term, the $W_{\varepsilon}$ distance between a distribution and itself is not 0, which can introduce biases when used to optimize generative models. To allow for effective optimization while retaining the computational advantages of entropic regularization, a common alternative is the Sinkhorn Divergence (Genevay et al., 2018), which subtracts the self-transport costs from $W_{\varepsilon}$:
@@ 3.2 Multi-dimensional Scaling
From (Gower, 1985), a matrix $D$ is a Euclidean Distance Matrix if and only if all of its values are non-negative, it is hollow (diagonal elements are 0) and the criterion matrix $F$ is positive-semidefinite (PSD):
@@
Where $J=I_{N}-\frac{\mathbf{1}_{N}\mathbf{1}_{N}^{\top}}{N}$ is the centering matrix and $1_{N}$ is a vector of length $N$ filled with ones.
Wasserstein distance is not Euclidean (see section 8.3 in (Peyré et al., 2019) and (Naor & Schechtman, 2007)), and there are many examples of cohorts of distributions (Figures S1,S2) wherein pairwise OT distance matrices do not fit the criteria in equation 4. Unlike the Euclidean scenario where classical MDS (cMDS) improves as embedding size grows, on non-Euclidean distance matrices the error behaves counter-intuitively and worsens as dimension increases (Sonthalia et al., 2021). This motivates us to seek a description of the theory underpinning the embeddings of non-Euclidean distances.
@@ 4 Wasserstein Wormhole
The complexity of modern OT solvers is quadratic with respect to the number of points in a point cloud; thus, it is only straightforward to compute pairwise Wasserstein distance for small cohorts of empirical distributions. Indeed, naive approaches will not even scale to standard datasets, such as the 12,311-point-cloud ModelNet40 benchmark (Wu et al., 2015).
  • Point clouds contain varying numbers of samples. $T(X)$ must be able to generalize to arbitrary cloud sizes while modeling relationships between the points within them.
@@
  Input: point clouds $\Omega=\{X_{i}\}^{N}_{i=1}$, initialized encoder $T$ and decoder $G$ networks, batch size $B$, learning rate $\varepsilon_{t}$
@@
By training the encoder network $T(X)$ (Figure 1, Algorithm 1), we can produce encodings which allow for OT-based analysis of point clouds in linear time, enabling methods such as clustering and visualization (Bachmann et al., 2022) on the entire cohort. However, the encoder alone does not provide a clear interpretation of these analyses.
Fig. 3: Wormhole on MNIST point clouds. a. Example of a point cloud produced by thresholding MNIST images. b. Retrieval of the Wasserstein distance in Wormhole embedding space. From a random sample of $128$ point clouds from the test set, we denote the correlation between the true pairwise OT distance and the Euclidean distance between their embeddings. The line $y=x$ is drawn in blue for reference. c. 2D UMAP visualization of Wormhole embeddings of the training and test set point clouds, which recapitulates ground-truth image digit without utilizing label information during training.
@@ 5 Embedding Non-Euclidean Distance Matrices
In the case where $D$ is Euclidean, by definition there exists a set of embeddings in dimension $d=N$ where the distance matrix is perfectly reconstructed and $\mathcal{L}=0$. Wormhole, however, is trained to minimize the stress relative to the Wasserstein distance matrix, which is generally not Euclidean. It is currently unclear to what extent $D$ can be represented with embeddings when it is not Euclidean (Sonthalia et al., 2021).
@@
This bound also appears in (Sonthalia et al., 2021), which analyzed pathologies of cMDS on non-Euclidean distance matrices.
@@
Algorithm 2 Theoretically Optimal Embedding of non-Euclidean Distance Matrices with Projected Gradient Descent (PGD)
  Input: Distance Matrix $D$, Initial Solution $D^{\prime}_{0}$, Learning Rate $\varepsilon_{t}$
@@
     $D^{*}_{t+1}=D^{\prime}_{t}-\varepsilon_{t}\nabla(\|D^{\prime}_{t}-D\|_{F}^{2})$GD Step}
@@
This new projection operator, along with the simple projections onto $C_{2}$ (hollow matrices, by setting the diagonal to 00) and $C_{3}$ (non-negative matrices, by thresholding), allows us to formulate a PGD (Boyd & Vandenberghe, 2004) scheme (Algorithm 2) to embed non-Euclidean distance matrices. This algorithm projects onto $C_{1}\cup C_{2}\cup C_{3}$ using Dykstra iterations (Boyle & Dykstra, 1986) and is guaranteed to find the optimum for equation 6, unlike other MDS approaches such as cMDS or SMACOF (De Leeuw & Mair, 2009). The PGD algorithm requires knowing the full pairwise distance matrix and is therefore infeasible in our setting of massive cohorts. However, for small cohorts, the theoretical assurances of PGD provide a useful benchmark for Wormhole.
Our analysis indicates that in real-world datasets, the Wormhole approximation is highly accurate and commensurate with theoretical values. Although this assessment is based on a small subsample of the full dataset, we assume that if random subsets of the larger dataset can be well embedded, the embedding of the entire dataset is reliable (Table 2). Full derivation of the theoretical bounds appear in Supplementary Sections 9 & 10.
Dataset Lower Upper PGD Wormhole cMDS
Simplex ($35$) $0.765$ $6.369$ $1.117$ $1.420$ $7.134$
Gaussian ($128$) $0.129$ $2.552$ $0.168$ $0.401$ $2.681$
MNIST ($256$) $0.042$ $0.616$ $0.058$ $0.100$ $0.657$
Table 2: Comparison between embedding algorithms and theoretical bounds on small datasets where full Wasserstein distance matrices are realizable. Values are the stress between the embedding and true distances and parentheses denote the size of each dataset. Classical MDS produces far from optimal errors, falling outside our bounds. Wormhole training is stochastic, yet it still converges between the error bounds and approaches the error obtained by the theoretically optimal PGD. Both PGD and Wormhole embeddings were well correlated ($\rho=0.99$) with OT distances.